OpenStudy (yttrium):

Find the 2014th digit after the decimal point of the decimal expansion of the repeating decimal 2/13.

OpenStudy (anonymous):

What do you have at your disposal? I have this trick with power series in mind, but I was thinking maybe finding a preliminary pattern by long division might be a good start.

OpenStudy (anonymous):

The pattern should repeat after \(n\) digits since it is a rational number. You just need to use \(\pmod n\)

OpenStudy (anonymous):

And hope that \(n\) isn't huge.

OpenStudy (yttrium):

It has a pattern!

OpenStudy (anonymous):

It's not, long division yields \(\dfrac{2}{13}=0.\overline{153846}\).

OpenStudy (yttrium):

Then we can easily solve for it

OpenStudy (anonymous):

\[ f(x) = f(x+6) = f(x+9k) \]

OpenStudy (yttrium):

2014/6 = 335 remainder 4.

OpenStudy (anonymous):

make that 6

OpenStudy (yttrium):

then the answer will be 8, right?

OpenStudy (anonymous):

I'm not up to date on the modular arithmetic, but I'll trust your judgment @wio :)

OpenStudy (anonymous):

Yeah, if you consider \(8\) to be the 4th digit, then \(8\) would be the 2014th digit as well.

OpenStudy (yttrium):

I'm sure it will yield with same answer.

OpenStudy (yttrium):

Okay. That's the easy part. How if the decimal is non repeating and non terminating? Is there any way to solve it?

OpenStudy (anonymous):

I think you just gave a description of an irrational number. A computer would be able to figure that out for you.

OpenStudy (anonymous):

Well, solving for a particular digit of a rational number? I dunno.

OpenStudy (anonymous):

irrational number^

OpenStudy (anonymous):

WolframAlpha has the capacity: http://www.wolframalpha.com/input/?i=2014th+digit+of+pi

OpenStudy (dan815):

is that izzy from digimon

OpenStudy (dan815):

i don't know if i got his name right

OpenStudy (anonymous):

@SithsAndGiggles Dunno if it actually calculates it out that far or calculates the specific digit.

OpenStudy (anonymous):

My input is incorrect, yeah, should be "2015th digit", but the function is still there.